OpenStudy (anonymous):

tanyas taking cookies to a bake sale. she already made 70 cookies on monday, and wants to make more on tuesday. her pan allows her to bake 18 cookies per batch. how many cookies will tanya have if she bakes 12 more batches y=70+18(x)

OpenStudy (anonymous):

where x=12

OpenStudy (anonymous):

x=12 y=286 y=70+18x y=70+18(12) y=70+216 y=286